# Search for gravitationally lensed interstellar transmissions

### v1

@@ V Discussion
Every star is destined to form an Einstein ring. This ring may form naturally with distant sources or through advanced technology, deliberately deployed to send us a message. The position of an Einstein ring around a star, as detailed in (10), depends on our distance to the star, the star’s mass, and the transmitter’s location. With this information, we can initiate our search for the laser transmissions enabled by gravitational lensing.
Through a synergistic effort, we will soon travel to the focal region of the solar gravitational lens (SGL), enabling direct high-resolution imaging and spectroscopy of exoplanets in our galactic neighborhood Helvajian et al. (2023); Friedman et al. (2024). However, before this, we must also survey nearby stars for potential signs of signals sent through gravitational lensing-enabled transmission links. Upon the mission’s arrival at the focal region of the SGL, we will have the capacity to observe the surface of exoplanets and initiate meaningful information exchanges with any inhabitants.
@@
We explored the case where a transmitter lies in the focal region of a nearby star, with the receiver positioned within our solar system. Although one must account for other light sources near the lens, their impact on the achievable SNRs is expected to be minimal. Given the high sensitivities, we consider the strategies in the search for transmitted signals using existing astronomical facilities or involving development of new facilities dedicated for this purpose.
Our findings reveal that microlensing events, corresponding to laser signals transmitted through nearby gravitational lenses, are transient but notably bright, making them detectable with current instruments. Given their brightness, these microlensing events can be detected and analyzed through a dedicated campaign of photometric observations. Deploying a network of astronomical facilities can enhance detection sensitivity. Consequently, we have demonstrated the feasibility of establishing interstellar power transmission links via gravitational lensing and also our technological readiness to receive those signal—a finding with profound implications in the search for interstellar power transmissions. Now is the right time to design, develop and initiate an appropriate search campaign.
@@ Acknowledgements.
This work was performed at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration.

### v3

@@ V Discussion
Every star is destined to form an Einstein ring. This ring may form naturally with distant sources or through advanced technology, deliberately deployed to send us a message. The position of an Einstein ring around a star, as detailed in (20), depends on our distance to the star, the star’s mass, and the transmitter’s location. Although we can not resolve the rings that may form around nearby stellar lenses by light transmitted to us from their focal regions, with this information, we can initiate our search for the laser transmissions enabled by gravitational lensing.
Through a synergistic effort, we will soon travel to the focal region of the solar gravitational lens (SGL), enabling direct high-resolution imaging and spectroscopy of exoplanets in our galactic neighborhood Helvajian et al. (2023); Friedman et al. (2024); Turyshev et al. (2023). However, before this, we must also survey nearby stars for potential signs of signals sent through gravitational lensing-enabled transmission links. Upon the mission’s arrival at the focal region of the SGL, we will have the capacity to observe the surface of exoplanets and initiate meaningful information exchanges.
In the early phase of this process, a network of astronomical ground- and space-based astronomical facilities may provide valuable spatial, spectral and temporal information as microlensing events unfold. This data may be processed with existing tools Turyshev and Toth (2023) to recover the properties of the signal and potentially the information encoded within.
@@
We explored the case where a transmitter is in the focal region of a nearby star, with the receiver positioned within our solar system. Although one must account for other light sources near the lens, their impact on the achievable SNRs is expected to be minimal. Given the high sensitivities, especially in the well-motivated signal-dominated transmission regime, we considered the strategies in the search for transmitted signals using existing astronomical facilities or involving development of new collaborative spatially dispersed facilities dedicated for this purpose.
Our findings reveal that microlensing events, corresponding to laser signals transmitted through nearby gravitational lenses, are transient, requiring good and deliberate pointing on the part of a transmitter but notably bright and spatially widened, making them detectable with the current generation of instruments. Given their brightness, these events can be detected and analyzed through a dedicated campaign of photometric and spectroscopic observations. Deploying a spatially disperse network of collaborative astronomical facilities can enhance detection sensitivity. It may also be used to learn about the transmitter and to determine if any message was transmitted. As a result, we have demonstrated the feasibility of establishing interstellar power transmission links via gravitational lensing, while also confirming our technological readiness to receive such signals. It’s time to develop and launch a search campaign.
Clearly, more work is still required, including: i) A comprehensive analysis of the propagation of annular Gaussian laser beams in the vicinity of gravitational lenses, as was done in Turyshev and Toth (2017) for plane waves. ii) A detailed exploration of the caustic patterns formed in the focal region of a gravitational lens by incident annular beams, similarly to Turyshev and Toth (2021b). iii) Consider spectral signals available with gravitational lensing, as in Turyshev and Toth (2022). iv) Numerical simulations to assess various transmission scenarios using incident Gaussian laser beams. v) Development of realistic SNR estimates using practical values for several key parameters, including detector quantum efficiency, optical throughput, spectroscopic sensitivity, instrumental noise, facility’s location, among technical capabilities essential for the search. vi) Formulating requirements for reliable signal detection by conducting a comprehensive exploration of the available set of parameters. vii) A systematic analysis of the capabilities of existing facilities and those under construction to support the search efforts. Research in these areas is ongoing and results, when available, will be reported in subsequent publications. The outcome of this work could have profound implications for the ongoing search for interstellar transmissions.
@@ Acknowledgements.
We acknowledge discussions with Jason T. Wright who kindly provided us with valuable comments and suggestions on various topics addressed in this document. This work was performed at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration.